OpenStudy (anonymous):

Which expression is equivalent to Which expression is equivalent to 5(c + b) + 4c – 2b?

OpenStudy (aravindg):

Okay first distribute 5 to c+b

OpenStudy (anonymous):

What?

OpenStudy (aravindg):

5(c+b)=?

OpenStudy (texaschic101):

Example : 2(x + y) = 2x + 2y Thats how you distribute through the parenthesis

OpenStudy (texaschic101):

you multiply the number outside of the parenthesis by everything inside the parenthesis

OpenStudy (anonymous):

Oh okay :P

OpenStudy (texaschic101):

5(c + b) = can you distribute the 5 through the parenthesis ?